OpenStudy (anonymous):

Solve this System... -x+2z=-9 -x-3y-4z=2 -3x-2y+2z=17

OpenStudy (anonymous):

Is it No Solution?

jimthompson5910 (jim_thompson5910):

what makes you say that

jimthompson5910 (jim_thompson5910):

Did you end up with a contradiction when you tried to solve this system?

OpenStudy (anonymous):

@butterflyprincess http://www.google.com/search?tbs=sbi:AMhZZitU0cfig31RKj2B3PWn8QM2KUc2Lm6CBtX76Rmh908jEYD_12jPmSd24WxjSNt3FhttCTtSxlGDeWbnDsKCrOnWp9qhKsoqXIXQ1fXzvbj4ueuVm2GQvFxzhgCARSP6kpP1tn_1wF4eJ3fLlgUdqfwaKOEjYeJnJn_1F_1nFKaYhxoOu3o1lp7aVgd6SEU_1WN69Uu8coQkD-x3TDB220NdqYJhusdwdaPEnprHW35O5lJ2IZ4UIwOL4E97qd_1aGy-q76pBBz5ECp8oTQ-tq0qPxAxwNVBxG_1JYaOcIeaFDJOmjk-bqA61NLR-Ow7HFdKi826sJsIwhJ0ZbERTQrdibjewVHrC9TqaQRB8n99vy_19EhQjL9asLfTct_1DRQ86ZSkCZB1TacrIfsyFwiRrncB09PD7dNPD2BstLbmKfjAydG2ZIi4Q1ZEC8IjEHgTEk7GIWSHXlZ655CIZCedcepfOZT2ha46rWI_1oiRmwochd_1U-jgq2LG4B0BH39VOlGMJGvcRMaR0xisdKm2UIRAgPGOmA77QM6AjYhtXNk_1dEBdJ5ei7NrzY1x9EiJFHdfAdwvfXm8m_1dHdZmGHTAMDQgHgOG84FxNaFFESyc-13-T33ldpmQZbLQK4RGhAfsLyID1aG4kzYwoJ3beKVnGzkAD9WxhMmWpUDggBwW-HaSIx4XR_1UrN4OhMp7TkEg_1iHowDGqWt6LRhyrPCChT4Y6z7zwiNY67A2aGUUmgmGjUTeN2gdvqHll8VGNUunkRZf_1tixEjz8QnbXqxeR0hrnXHJC86vhgPRV_1hSzMH3qnbTwQ10cMh8IhdyoCaDzGiC4g4rAF0SU3aObGkm6GDdCHH-gGr5K_1fDSDhrqGlEeNrAg2ITbBhXVn0sD8bBwXye_1ye7uuo7sqhOre1S9JA2lOoKg45fhWvyMHkUg9d9FvULfZbLoSvSL9E7CZjoaxbFBt-k9a8VHF6ZUNSliG8A_1yWj4esjACyJcwVvkreAXmGMxonnVJIG1QH-8L8heMn-sYTdRpBDr24vKPNAG5sBN-4S1vhBmPocUwg_1EhmMgXoaln3ASuWn8M0UTCyBbW7BW2DfJF1aSNLhG-g1UT-jNDJt3RtTBRCuWAYgkf0GTn2xNDYzZbe0ira7LXiuB91Sb_12VhHPHtxvyzQhwJUGbHWL6Mdma01dbyao_10or1sZA68v3B0wk6e-of8mOR56RlXG-MXJanChX2_1qdmp-nECg4E0j2_1jVKWQ8v4Ld0vSlz5X-35zPMP7ZDYxh2Rv2dTZF_1MQMunsG9tfG24E02fxSKyom7jg6h1xubVO28XCWBSDMJ0G62OdFI6GovXQ2CX3UdMwqQWuNG5hduxZbx2-CLQeRmwuZg8i5B_1P3nZhyTv78BU3fOxKqQ&=Search&num=10&hl=en&bih=750&biw=1069

OpenStudy (anonymous):

I ended up with like -21y-42z=77 21y+42z=33 it cancels out

OpenStudy (anonymous):

you could be charged with identity theft.. i would change it...

jimthompson5910 (jim_thompson5910):

here's how I would do it -x+2z=-9 -x = -2z - 9 x = 2z + 9 ---------------------------- -x-3y-4z=2 -(2z + 9) - 3y - 4z = 2 -2z - 9 - 3y - 4z = 2 -3y - 6z - 9 = 2 -3y - 6z = 2+9 -3y - 6z = 11 ------------------- -3x-2y+2z=17 -3(2z + 9)-2y+2z=17 -6z -27 - 2y+2z=17 -2y - 4z - 27 = 17 -2y - 4z = 17+27 -2y - 4z = 44 --------------------------------- -3y - 6z = 11 -2y - 4z = 44 -2(-3y - 6z) = 2*11 3*(-2y - 4z) = 3*44 6y + 12z = 22 -6y - 12z = 132 ---------------- 0y + 0z = 154 0 = 154 ... which is a contradiction So you are 100% correct. There are no solutions and this system is inconsistent.